# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

### v1

@@ 4.1 Overview of CARE Training Platform
Figure 5: CARE’s practice and feedback model visualized in a diagram and web screenshot. In CARE, counselors practice with an LLM-simulated patient and receive feedback on each of their responses. The feedback model labels whether a response has strengths or constructive feedback areas. Responses with constructive feedback explain what the goal should be at this point in the conversation; what a helper could improve to better align with this goal; and how they could respond differently via an alternative response.
@@
We developed CARE as a web platform for novice counselors to train in text-based counseling skills enabled by LLMs. Trainees practice simulated conversations with AI patients and review feedback on their responses generated by an AI mentor. CARE builds on top of successes from previous research in co-designing with mental health experts to improve the realism of LLM-simulated patients [50, 43, 14, 15] and finetuning domain-specific LLMs with therapeutic knowledge capable of generating feedback and alternative responses for text-based peer counseling conversations [51, 22, 52, 53, 54].
CARE allows novice counselors to develop their counseling skills in a text-based format by practicing with AI-simulated patients and receiving feedback on their responses (see Fig. 5).
@@
Consider Aki, a novice peer counselor who wants to use CARE to experience hands-on training using counseling skills that they have recently read about. In CARE, Aki can practice with an AI patient of their choice from a library of patient scenarios. Aki starts a practice chat with an AI patient with the background of a 35-year-old male veteran seeking to reconnect with his children but facing legal barriers and parental gatekeeping.
@@
Each practice scenario provides limited background information about the AI patient and their presenting problem (e.g., ”Young adult with family issues: low mood and self-esteem”). This intentional limitation requires counselors to simultaneously learn more about the patient’s situation while demonstrating empathy and support.
@@
Aki starts the conversation by greeting the AI patient. The AI patient shares their frustration with feeling everyone is against him, and hopes to find ways get his kids back and overcome the difficulty his judgmental parents. Aki composes a reply, and the conversation continues in this turn-by-turn manner.
Users can view feedback on their therapeutic responses at any point during or after the session, offering flexibility to either review intermittently throughout the practice or comprehensively after completion.
4.2 Implementation details for CARE Training Platform
CARE was built as a web application with a Python Flask back-end and React Javascript front-end which online participants could access and interact with. LLM-simulated patients in CARE were powered by the GPT-4o API with a custom prompt defining the patient’s background scenario and behavioral principles [15]. The prompts for the three AI patients used for the pre-intervention assessment, intervention, and post-intervention are provided in the Supplementary Materials A.1. The LLM feedback system was powered by a Llama-2 13B parameter model finetuned on an expert-annotated feedback dataset [22].
4.3 Study Procedures
Between August and October 2024, we conducted an online lab study with (N=94) novice counselors. Given our aim to train novices lacking formal counseling training, we recruited participants who expressed an interest in practicing helping skills yet had minimal prior education or practical experience. Specifically, we used the Prolific platform to screen for participants who (1) lacked extensive formal professional or graduate training in clinical psychology, counseling, social work, or nursing (excluding those who were pursuing or had completed master’s or doctorate degrees); and (2) had less than one year of counseling-related experience (e.g., peer support or crisis counseling). In our final experiment, we recruited N=94 participants. Additionally, during the pilot phase, we recruited another N=14 participants who were allowed broader educational and hands-on experiences for protocol refinement, but we report qualitative and quantitative results exclusively from the final set of 94 participants.
Each participant completed a 75-90 minute video-call session following the structure outlined in Fig. 1. After receiving a static 5-minute tutorial on core counseling skills (see Fig. 6), all participants completed a 10-minute pre-intervention assessment chat with the first AI patient. For the 20-minute training intervention, participants were randomly assigned to either practice with an AI patient only or practice with the option to receive feedback on all of their sent responses. Finally, all participants completed a 10-minute post-intervention assessment chat with a third AI patient. Participants answered self-efficacy items after the pre-intervention and post-intervention chats, and provided qualitative self-reflections after the training intervention chat and post-intervention chat. All participants were paid \$15/hour. To protect participant privacy, we instructed them to join the video-call using their ProlificID as the display name, turn off their video-camera, and share their screen. Participants consented for the video call to be recorded for our own analysis and their CARE usage data to be used and shared for the purposes of this and future research after de-anonymization. This study protocol was approved by our institution’s IRB.
4.4 Assessing Changes in Skill Use through LLM Classifiers
We investigate whether counselors employ higher-quality counseling behaviors after LLM-based training and between training conditions by leveraging NLP methods. This analysis is motivated by the need to quantify changes in counseling skill use at scale across multiple participant sessions. Our approach follows four key steps: (1) fine-tuning and validating LLM-based classifiers to identify skill behaviors, (2) selecting a final set of classifiers based on performance metrics and theoretical priority, (3) testing changes in skill use across time and between intervention groups, and (4) analyzing the relationship between exposure to good alternative patterns in AI feedback and subsequent skill use.
4.4.1 Fine-tuning and Validating LLM-based Classifiers
We developed LLM-based binary classifiers that allows us to label the skill use within a transcript. For example, one of our fine-tuned classifier could determine which utterances in a transcript had strong uses of Questions during that stage of the transcript. To finetune and evaluate these classifiers, we transformed a previously published expert-annotated, feedback dataset that had strengths and areas of improvements for 8 skills [22] into a 16-class binary classification format (8 skills × 2 categories each—strong uses and areas needing improvement) [^1] The binary classification feedback dataset can be accessed at https://huggingface.co/datasets/youralien/feedback_qesconv_16wayclassification. Additionally, we used a subset of transcripts from this randomized study which were further annotated by counseling domain-experts.
@@ 4.4.1 Fine-tuning and Validating LLM-based Classifiers
Three experts were recruited with relevant backgrounds including practicing clinical psychologist, licensed marriage family therapist, former director and supervisor of a crisis agency. Each of the experts annotated 10 participants’ study transcripts (5 from the practice-only group; 5 from the practice-and-feedback group), totaling 370 counselor utterances. Two rounds of annotation occurred: after collecting a first annotation pass and identifying data points with disagreements, we showed each of the experts the others’ annotations and had them re-annotate and provide rationales for their decision. We display pairwise agreement results averaged across all pairs in Table 2. Note that while we originally explored other annotation agreement metrics such as Cohen’s kappa, the severe class imbalance of our annotations made these metrics less relevant in our case. The gold-standard CARE 10% expert-annotated sample consists of labels that result from a majority vote across these three experts [^2] This expert-annotated data sample can be found at https://huggingface.co/datasets/youralien/CARE_10percent_16wayclassification.
We finetuned RoBERTa-large binary classifiers using a 95% split of the transformed FeedbackQESConv dataset, and did hyperparameter tuning on 5% of FeedbackESConv and the CARE 10% sample. The performance of the our LLM-based classifiers are shown in Table 2. The highest performing classifiers ($F1>0.5$) became candidates for our automatic behavioral assessments, which we further down-selected as described below.
@@
Table 2: Left columns show pairwise annotator-agreement averaged across 3 domain-experts for the CARE 10% sample (n=370). Right columns show performance of the best RoBERTa-large classification models after hyperparameter tuning on our validation dataset, CARE 10% sample (n=370) + FeedbackQESConv 5% sample (n=409). Session Management and Professionalism were excluded from finetuning due to infrequent occurrence.
4.4.2 Down-selecting a Final Set of Classifiers
From the initial set of 16 binary classifiers, we applied both methodological and theoretical criteria to select our final set for analysis. First, we established a minimum performance threshold of $F1>0.5$ to ensure reliable classification. This criterion yielded seven candidate classifiers: strong uses of Empathy, Reflections, Questions, and Validation, as well as both strong uses and areas needing improvement for Suggestions.
To maintain statistical power while controlling for multiple comparisons, we further narrowed our focus to four key classifiers: strong uses of Empathy, Reflections, and Questions, plus areas needing improvement for Suggestions. This selection was guided by three considerations: (1) the need to limit the number of statistical comparisons to avoid diluting significance across too many tests, (2) focusing on classifiers with the strongest performance metrics, and (3) selecting skills that represent core competencies in client-centered frameworks and are frequently used in counseling sessions.
Self-disclosure was excluded from our analysis due to its infrequency in our dataset, while Validation, though conceptually related to Empathy and mentioned frequently in qualitative data, showed more limited classifier performance and was therefore reserved for secondary analyses.
4.4.3 Testing Changes across Time and between Groups
Using our selected classifiers, we examined how skill usage changes from the pre-intervention chat to the post-intervention chat. For each session transcript, we computed an aggregate score defined as the proportion of utterances with strong uses of a skill (e.g., $b=N_{strengths}/N_{total}$) or with areas needing improvement (e.g., $b=N_{improvement}/N_{total}$).
To statistically test for changes between the pre-intervention and post-intervention chats ($b_{0},b_{1}$), we used paired t-tests and computed Cohen’s $d$ effect sizes. To test for significant differences between P and P+F groups ($b_{1}^{P}-b_{0}^{P}$ vs. $b_{1}^{PF}-b_{0}^{PF}$), we used non-paired t-tests.
In total, we conducted 12 planned t-tests for analyzing changes in behaviors: three skills (Empathy, Reflections, Questions) with strong uses and one skill (Suggestions) with areas needing improvement, each analyzed for both within-group changes and between-group differences in those changes.
4.4.4 Exposure to Good Alternatives in AI Feedback
To better understand AI feedback’s role in skill development, we analyzed whether the level of exposure to specific feedback during practice affected post-intervention performance. We quantified the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length.
To test whether greater exposure to GAP for a skill led to stronger use of that skill, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$). This model allowed us to isolate the effect of feedback exposure while accounting for baseline performance differences.
4.5 Analyzing Confidence and the Calibration of Self-Assessments using Counselor Self-Efficacy
Counselor self-efficacy represents a critical dimension in counselor development, defined as one’s “beliefs or judgments about their ability to effectively counsel a client” [55]. While high self-efficacy can promote persistence and resilience when facing challenges [56], miscalibrated self-assessments—particularly overconfidence relative to actual ability—can lead to suboptimal or potentially harmful counseling behaviors [57, 58, 59].
This calibration challenge has prompted research into potential remedies. Prior studies suggest that increased practice experience may help individuals better align their self-assessments with actual performance, though results remain inconsistent [60]. Similarly, supervisor feedback has been proposed as a mechanism to address self-assessment miscalibration [61], although evidence from other educational contexts shows mixed effectiveness [62]. Our methodology specifically examines whether practice alone or practice with structured feedback differentially impacts both raw self-efficacy and its calibration with actual performance.
4.5.1 Self-Efficacy Measurement Instruments
@@
To measure counselor self-efficacy, we employed the Counselor Activity Self-Efficacy Scale (CASES) [37], specifically utilizing a revised subset of items targeting basic counseling skills (CASES-R) [63, 36]. The CASES-R established a three-factor structure to assess counselors’ confidence in performing key therapeutic functions: Exploration and Insight Skills (EIS), Action Skills (AS), and Session Management Skills (SM).
@@
Participants completed the CASES-R immediately following both pre-intervention and post-intervention AI patient interactions. All items were administered using an 8-point Likert scale (0 = no confidence, 7 = complete confidence). During factor analysis, we discovered that among the five original EIS skills, self-disclosure did not load on the same factor as the other items. Consequently, we consolidated the EIS dimension to include only four Exploration Skills: Reflections of Feelings, Restatements, Open Questions, and Listening.

### v2

@@ 3. CARE Training System
Figure 2. CARE’s practice and feedback model visualized in a web screenshot. In CARE, counselors practice with an LLM-simulated patient and receive feedback on each of their responses. The feedback model labels whether a response has strengths or constructive feedback areas. Responses with constructive feedback explain what the goal should be at this point in the conversation; what a helper could improve to better align with this goal; and how they could respond differently via an alternative response.
The figure shows a screenshot of CARE’s practice and feedback interface. On the left, a text-based chat unfolds between an AI patient (depicted with a client avatar) and a counselor (depicted with a clinician avatar). The AI patient expresses frustration about his parents’ lack of trust and his struggles to prove he is a good father. The counselor responds twice: first with an empathetic statement that acknowledges the difficulty, and second with a suggestion-oriented prompt about showing efforts to change. On the right, CARE’s feedback system evaluates these responses. The first counselor response is labeled a Strength, tagged as “Empathy,” with positive feedback saying “Good response!” The second counselor’s response is flagged as a Feedback Area, tagged under “Suggestions” and “Questions.” The system provides an Alternative Response that rephrases the counselor’s statement to be more client-centered and exploratory, asking about the patient’s efforts rather than giving advice. Additional Feedback text explains that the goal at this point is to explore the seeker’s feelings and concerns rather than provide direct solutions, encouraging open-ended questioning.
We developed CARE as a web platform for novice counselors to train in text-based counseling skills enabled by LLMs. The system integrates two core components: (1) LLM-simulated patients that provide realistic, text-based practice conversations, and (2) LLM-generated feedback that evaluates counselor responses against established skill frameworks and suggests improvements. Together, these features enable scalable, authentic training experiences that complement traditional, resource-intensive approaches such as role-play and supervision.
@@
CARE builds on top of successes from previous research in co-designing with mental health experts to improve the realism of LLM-simulated patients (Chen et al., 2023; Lin et al., 2024; Wang et al., 2024; Louie et al., 2024), using fine-tuned domain-specific LLMs trained on therapeutic knowledge capable of generating feedback and alternative responses for text-based peer counseling conversations (Sharma et al., 2021; Chaszczewicz et al., 2024; Shen et al., 2022; Min et al., 2022, 2023). Importantly, CARE was designed not only to identify whether a skill is used but also how well it is used, distinguishing, for example, between a reflection that captures a client’s core concern and one that misses the emotional nuance. CARE allows novice counselors to develop their counseling skills in a text-based format by practicing with AI-simulated patients and receiving feedback on their responses (see Fig. 2).
@@
Consider Aki, a novice peer counselor who wants to use CARE to experience hands-on training using counseling skills that they have recently read about. In CARE, Aki can practice with an AI patient of their choice from a library of patient scenarios. Aki initiates a practice chat with an AI patient, a 35-year-old male veteran who is seeking to reconnect with his children but is facing legal barriers and parental gatekeeping.
@@
Each practice scenario provides limited background information about the AI patient and their presenting problem (e.g., ”Young adult with family issues: low mood and self-esteem”). This intentional limitation requires counselors to simultaneously learn more about the patient’s situation while demonstrating empathy and support. Patients are designed to exhibit realistic challenges, including resistance, ambivalence, or vague disclosures, drawing on behavioral principles elicited from expert counselors (Louie et al., 2024).
@@
Aki starts the conversation by greeting the AI patient. The AI patient expresses frustration about feeling that everyone is against them, hoping to find ways to reunite with their kids and overcome the challenges posed by their judgmental parents. Aki composes a reply, and the conversation continues in this turn-by-turn manner.
After completing the conversation, Aki reviews AI-generated feedback. CARE highlights strengths such as asking open-ended questions, but also flags missed opportunities for empathy, offering alternative phrasings, and a rationale for why they may better support the client.
Unlike real-time AI ”co-pilot” systems, which may proactively suggest responses, CARE provides feedback only after the user has sent their response in the simulated dialogue. A user can view feedback on their therapeutic responses at any point. This offers flexibility to review feedback intermittently throughout the practice or comprehensively after completion. This design choice mirrors human supervision: it preserves the learner’s agency during the conversation while supporting reflection afterward. Feedback targets core microskills, empathy, reflections, questions, validation, and suggestions, based on established counseling frameworks (Nelson-Jones, 2013; Miller et al., 2003).
3.1. Implementation details for CARE Training Platform
CARE was built as a web application with a Python Flask back-end and React JavaScript front-end, accessible through any standard browser. The two core components of CARE are described below.
LLM-simulated patients. CARE implements an existing method for simulated patients that co-designed patient prompts with expert-counselors, each specifying (a) a challenging patient scenario that they had encountered in past clinical and text-based counseling settings, including demographic background presenting issues or symptoms, etc. and (b) Constitutional AI principles elicited from expert counselors for defining authentic patient behaviors (Louie et al., 2024). This simulation method uses the OpenAI GPT-4o API to role-play patient scenarios and behaviors due to its strong ability to maintain role consistency and instruction-follow expert-defined principles. These principles instructed the simulated patients to display realistic challenges such as resistance to suggestions ("Respond to encouraging words with hesitation, doubting their significance"), low awareness ("Don’t be so self-aware or good at recognizing your own problems"), or minimal disclosure ("Use more colloquial language and express reluctance to open up"). Simulated patients were designed and validated for 10-20 minute text-based counseling conversations, allowing participants to both explore the presenting problem and practice multiple client-centered microskills. CARE implements the top-rated patients from Louie et al. (2024)’s study which were judged by third-party counselors as being the most ready to be used as a training partner ($\geq 6$ average score on a 7-point Likert-scale). All prompt templates are available in Supplementary Materials A.1.
The three AI patients used in our experiment were created by experienced counselors from the USA with extensive backgrounds in mental health support (Louie et al., 2024). Each patient was deliberately varied to expose trainees to diverse yet comparably challenging scenarios:
  • Patient 1 (35-year-old American male experiencing holiday loneliness): Created by an experienced peer counselor from an online counseling platform. Holiday loneliness is common there and often stems from family estrangement. In this case, the patient’s isolation intensifies during winter holidays when family gatherings occur.
  • Patient 2 (35-year-old male veteran in court-mandated therapy seeking to reconnect with children): Created by a Licensed Marriage and Family Therapist (LMFT), a white woman aged 30-40, based on her clinical experience with veterans. This scenario reflects a patient frustrated by limited access to his children due to parental and legal issues stemming from substance abuse.
  • Patient 3 (Young adult with family issues, low mood, and self-esteem concerns): Created by a US-basd clinical psychology doctoral student. This adolescent patient faced self-esteem issues stemming from family dynamics, where her parents favored a sibling, leading to symptoms of anhedonia and depression, with a diminished ability to enjoy previously pleasurable activities.
LLM-generated feedback. CARE integrates an existing method for generating counseling feedback that fine-tunes and self-improves the Llama-2 13B parameter model using an expert-annotated feedback dataset of peer counseling transcripts  (Chaszczewicz et al., 2024). We selected this existing method because it provides faithful generation of counseling feedback grounded in the content and style of feedback given by psychotherapy supervisors. An additional key benefit, as argued by Chaszczewicz et al. (2024), is that a fine-tuned open-weight model can operate on therapy data in a controlled, private environment rather than relying on external API services. This model generates feedback at multi-levels: (1) classify the trainee response against eight microskills (empathy, reflections, questions, validation, suggestions, session management, professionalism, and self-disclosure), (2) assess quality by highlighting strengths and areas needing improvement, and (3) generate alternative responses and explanatory rationales, enabling trainees to compare their choices against more client-centered approaches. This post-practice feedback design mirrors human supervision: it preserves agency during the conversation while supporting reflection and skill refinement afterward.
4. Randomized Pre-Post Study
The core goal of CARE is to upskill novice counselors through LLM-simulated training. In a randomized experiment, we investigated how CARE’s core components–practicing with LLM-simulated patients and receiving AI-generated feedback on their responses–are important for participants’ skill development. We conceptualize skill development holistically, encompassing three complementary dimensions: (1) behavioral performance, where a trainee is judged on their appropriate use of counseling skills in a representative scenario or conversation; (2) counseling self-efficacy, defined as a trainee’s self-assessments of their own abilities; and (3) therapeutic intentions, or the goals that participants form in-session, which should be adherent with evidence-based procedures.
The experiment evaluated how skill development changed over time when participants were assigned to two variants of LLM-based counseling training: practicing with LLM-simulated patients alone (Group P); or practicing with simulated patients while also receiving LLM-generated feedback (Group P+F). In contrast to other work on simulation in clinical education (Shin et al., 2015), our primary interest was in understanding what aspects of LLM-simulated training could promote skill development, as part of a broader research agenda to iteratively design and improve AI-based counseor training. Given this, we did not include a ”no AI” control group.
Beyond the skill development outcomes, we also conduct a mixed-methods investigation of participants’ experience of CARE’s LLM-based components and their perceived value of such training experience, since user perceptions shape adoption in training contexts. Together, our experiment sought to answer six research questions covering skill development and training experience with CARE from quantitative and qualitative perspectives.
  • RQ1: How does CARE’s LLM-simulated practice and feedback affect participants’ behavioral performance?
  • RQ2: How does CARE’s LLM-simulated practice and feedback affect participants’ self-efficacy?
  • RQ3: How does CARE’s LLM-simulated practice and feedback affect novice counselors’ therapeutic intentions?
  • RQ4: What are participants’ quantitative experience of CARE’s LLM-simulated practice and feedback?
  • RQ5: What are participants’ qualitative experience receiving feedback on their responses from CARE?
  • RQ6: What are participants’ qualitative experience practicing with CARE’s LLM patients?
4.1. Participants
We recruited $N=94$ novice counselors on the Prolific platform using specific filtering criteria to select US and UK participants with some interest in the field but limited access to formal training. Eligible participants were required to have (1) an educational background in psychology, counseling, social work, or nursing, with educational attainment limited to those who had completed at most a bachelor’s degree or were currently pursuing a master’s degree, and (2) less than one year of counseling-related experience (e.g., peer support or crisis counseling volunteering). Prolific participants were paid \$15/hour. We conducted sessions with 108 participants from Prolific. The first 14 were part of a pilot that refined our recruiting criteria and protocol (e.g., excluding counselors with graduate degrees). Our analyses use the final 94 participants, though we reference participants by their original identifiers. For the final participant pool, 68% were located in the United States and 32% in the United Kingdom. The sample was predominantly female (68%), with 31% male participants and 1% preferring not to disclose gender. The median age was 29 years (IQR: 23-39). Regarding ethnicity, 49.5% of participants identified as White, 16.2% as Black, 15.3% as Multiracial, 13.5% as Asian, and 5.4% as Other. Participants’ primary fields of study included psychology (66%), social work (24%), nursing (16%), and counseling (10%), with participants able to select multiple areas. In terms of educational attainment, 22.4% had no formal education in these fields, 50.6% were currently pursuing undergraduate degrees, 12.9% had completed only bachelor’s degrees in relevant fields, and 14.1% were pursuing master’s degrees.
To protect participant identities, our IRB-approved protocol instructed participants to use their Prolific email address, turn off cameras, and change their Zoom display name to their ProlificID. During recruitment and consent, we warned participants of the potentially stressful nature of simulated patient situations and ensured scenarios avoided especially sensitive topics such as suicidal ideation.
4.2. Power Analysis
To determine the appropriate sample size for our randomized pre-post study, we conducted a power analysis targeting a medium effect size with adequate statistical power. Our analysis was based on a repeated-measures design comparing pre-intervention and post-intervention outcomes between two groups (practice-only vs. practice-with-feedback). We selected an effect size of $d=0.4$ as our target, representing a conservative estimate for behavioral skill improvements. This choice was informed by previous research on social skills training interventions, where studies examining changes in behavioral performance have reported medium to medium-high effects ranging from $d=0.5$ to $d=0.6$ (Lin et al., 2024). Using power analysis calculations for two-sample, repeated-measures designs with $\alpha=0.05$ and $\beta=0.8$ (80% power) in the R statistical analysis software, our calculations indicated that $N=94$ participants would provide sufficient statistical power to detect our target effect size.
4.3. Study Setup
The study flow is illustrated in Figure 1. The 75-minute study was conducted over Zoom. Participants first read a 5-minute tutorial refreshing foundational counseling skills, then completed a timed 10-minute pre-intervention chat with the first AI patient. For the 20-minute main intervention, we randomized participants into two groups: (1) Group P: Practice with an LLM-simulated patient without AI feedback, or (2) Group P+F: Practice with an LLM-simulated patient with AI feedback. Group P+F participants could review AI feedback on their responses at any time. The experimenter provided verbal reminders to check feedback at 5 minutes and to review remaining feedback at 15 minutes. Participants then completed a 10-minute chat with the third AI patient. Surveys were administered after each chat period. Upon completing the post-intervention chat and self-efficacy assessment, participants shared their experience with the CARE training tool via survey and semi-structured interview. Group P participants received 5 additional minutes to interact with AI feedback on their post-intervention chat before sharing perceptions. Since this occurred after the skill acquisition experiment, it does not interfere with training effectiveness results (RQ1-3) but allows us to ask all 94 participants their perceptions of both AI patients and AI feedback in CARE (RQ4).
4.4. Measures
To understand whether simulated practice alone (P) and practice with feedback (P+F) can upskill novice counselors, we integrate evidence from three sources of data: automatic assessments of behavioral performance (RQ1), participants’ assessments of their self-efficacy (RQ2), and qualitative self-reflections about their therapeutic intentions (RQ3). Following the post-intervention, we conducted a final survey and semi-structured interview with participants to understand their perceptions of the CARE system and its features (RQ4).
4.4.1. RQ1. Automatic Assessment of Behavioral Performance
We assess whether counselors employ higher-quality counseling behaviors in transcripts by leveraging NLP methods. This automatic assessment is motivated by the need to quantify changes in counseling skill use at scale across multiple participant sessions. Our automatic assessment approach requires (1) fine-tuning and validating LLM-based classifiers to identify skill behaviors, and (2) selecting a final set of classifiers based on statistical-testing considerations, performance metrics, and theoretical priority. In the following paragraphs, we explain both of these steps in more detail. Ultimately, we assessed behaviors of skills used for the exploration stage (strong uses in empathy, reflections, questions) and action stage (suggestions needing improvement) of Hill’s Helping Skills framework (Nelson-Jones, 2013); see Table 2 for definitions.
Fine-tuning and Validating LLM-based Classifiers.
@@ 4.4. Measures
We developed LLM-based binary classifiers to label skill use within transcripts. For example, one classifier determines which utterances showed strong use of Questions. To finetune and evaluate these classifiers, we transformed a previously published expert-annotated feedback dataset (Chaszczewicz et al., 2024) into 16-class binary classification format (8 skills × 2 categories: strong uses and areas needing improvement). [^1] The binary classification dataset for FeedbackESConv can be accessed at https://huggingface.co/datasets/youralien/feedback_qesconv_16wayclassification Additionally, we used a subset of transcripts from this study, annotated by three counseling domain-experts: a practicing clinical psychologist, licensed marriage family therapist, and former director and supervisor of a crisis agency. Each expert annotated 10 participants’ transcripts (5 from each group), totaling 370 counselor utterances. After an initial pass, we showed experts each other’s annotations for disagreement points and had them re-annotate with rationales. Table 1 shows pairwise agreement results averaged across all pairs. While we initially explored metrics like Cohen’s kappa, severe class imbalance made them less relevant. The CARE expert-annotated sample (10% of participants’ transcripts) consists of majority-vote labels across the three experts. [^2] This expert-annotated data sample can be found at https://huggingface.co/datasets/youralien/CARE_10percent_16wayclassification
We finetuned RoBERTa-large binary classifiers using FeedbackQESConv, a dataset of transcripts from emotional support conversations between peer counselors on a crowdsourcing platform annotated with multi-level counseling feedback (Chaszczewicz et al., 2024), allocating 95% of this data for training. For hyperparameter tuning, we used a validation set comprising 5% of the FeedbackQESConv dataset (n=409) combined with our CARE expert-annotated sample (n=370). The performance of our LLM-based classifier candidates varied by skill, as shown in Table 1, motivating us to down-select a final set of classifiers for our planned analyses.
@@
Table 1. Annotator agreement columns show pairwise agreement averaged across 3 domain-experts for the CARE expert-annotated sample (n=370). Classifier performance columns show performance of the best RoBERTa-large classification models after hyperparameter tuning on our validation dataset, CARE expert-annotated sample (n=370) + FeedbackQESConv 5% sample (n=409). Session Management and Professionalism were excluded from finetuning due to infrequent occurrence.
Down-selecting a Final Set of Classifiers
From the initial set of 16 binary classifiers, we narrowed our focus to four key classifiers: strong uses of Empathy, Reflections, and Questions, and inappropriate uses of Suggestions. This selection was guided by selecting (1) the highest performing classifiers based on F1 scores (2) fewer classifiers for statistical concerns since our analyses would control for a false discovery rate based on number of skill hypotheses tested; and (3) theoretical relevance of skills most frequently emphasized in client-centered counseling textbooks and used during training with CARE. The final four skill classifiers have F1 performance scores between 0.56 and 0.77. Detailed selection criteria and rationale are provided in Appendix A.2.
4.4.2. RQ2. Counseling Self-Efficacy
@@
To measure counselor self-efficacy, we employed the Counselor Activity Self-Efficacy Scale (CASES) (Lent et al., 2003), specifically utilizing a revised subset of items targeting basic counseling skills (CASES-R) (Hahn et al., 2021; Hunsmann et al., 2024). The CASES-R established a three-factor structure to assess counselors’ confidence in performing key therapeutic functions: Exploration and Insight Skills, Action Skills, and Session Management Skills.
@@
Participants completed the CASES-R immediately following both pre-intervention and post-intervention AI patient interactions. All items were administered using an 8-point Likert scale (0 = no confidence, 7 = complete confidence). During factor analysis, we discovered that among the five original Exploration and Insight Skills, self-disclosure did not load on the same factor as the other items. Consequently, we consolidated the Exploration and Insight Skills dimension to include only four Exploration Skills: Reflections of Feelings, Restatements, Open Questions, and Listening.